Calculate the total amount of coupons of a bond that has a nominal value of $ 500, pays monthly interest at the rate of 15% annual capitalizable monthly and its redemption will be 5 years.
Answer #1

Total coupon paid over the life of the bond=Monthly coupon*12*t=500*15%/12*12*5=$375.00000